TIP: Click on subject to list as thread! ANSI
echo: sb-nasa_news
to: All
from: Hugh S. Gregory
date: 2003-03-07 22:56:00
subject: 2\12 Pt-1 ISS On Orbit Status 12-02-2003

This Echo is READ ONLY !   NO Un-Authorized Messages Please!
 ~~~~~~~~~~~~~~~~~~~~~~~~   ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
2\12 ISS On-Orbit Status 12-02-2003
Part 1 of 2

ISS On-Orbit Status 2/12/03

All ISS systems continue to function nominally, except as noted 
previously or below. Day 81 in space for Expedition 6.

After wake-up (1:00am EST) and before breakfast, FE-1 Nikolai Budarin
concluded his current MBI-8 "Profilaktika" (countermeasures) fitness
session, today conducting the blood test part, measuring lactate and
creatine kinase levels with the AccuSport equipment, and the physical
exercise on the TVIS treadmill. [The TVIS test is identical to the 
MO-3 test performed on the treadmill in idling mode with free choice 
of speeds within a certain specified range. In addition to the nominal 
test procedure, it calls for the use of the TEEM-100M gas analyzer 
during the test, the blood measurements, and subjective evaluation of 
physical exertion levels during the test. Gas analyzer and ECG 
(electrocardiograph) data were transferred to Laptop 3 and prepared 
for later downlink via Regul-Packet comm.]

Later, Budarin conducted a checkout of the Russian GAMMA-1M blood 
pressure (AD) monitor, by testing the AD cuff, with its cuff hose and 
cable, on a simulated arm (a rounded container).

CDR Ken Bowersox and Budarin continued unloading and transferring 
Progress 10P cargo, with IMS record-keeping.

FE-2/SO Don Pettit set up the acoustic dosimeters for static noise
measurements. Using the MEC (medical equipment computer) with special 
"EARQ" software, all crewmembers later recorded audiograms measured by 
their individual portable dosimeters for several hours.

Nikolai Budarin completed removal and replacement of two components of 
the SM BITS2-12 on-board measurement telemetry system, viz., an old 
memory unit (ZU1-A) and the BP11-A on-board transmitter. During the 
R&R, a number of Russian subsystems were temporarily disabled, such as 
the SKV-1 air conditioner, Vozdukh control algorithm, BMP 
micropurifier, Elektron, and the BITS itself. [The ZU memory device, 
one of four, uses 200m of magnetic tape to store and reproduce data. 
The BP11-A transmitter, one of four, is used to transmit telemetry 
data from the SM to the ground.]

Budarin also checked the gas analyzer in the SM, to verify its 
readings, then tagging up with ground specialists via S-band.

Later, Nikolai, working at the SM Central Post, continued and 
completed the battery cycling of the three new laptops delivered by 
Progress 10P.

Bowersox performed troubleshooting on EXPRESS Rack 2 (ER2). [He 
focused on two main approaches: (1) re-cabling the ER2 laptop back to 
nominal Ethernet configuration with a clean rebooted RIC (rack 
interface controller), then trying to re-establish Ethernet Comm. If 
successful, the ground was to continue with the loading of the laptop 
with the new software. (2) If not successful, ground was to continue 
with the second troubleshooting activity, which uses the ER3 laptop 
and Ethernet cable on ER2.]

Bowersox and Pettit conducted a refresh review of the DOUG (dynamic
operational ubiquitous graphics) software with new files for 
tomorrow's SSRMS/Robotics operations. [DOUG is a sophisticated 
software program that provides a graphic birdseye-view graphical image 
of the SSRMS (space station remote manipulator system), showing its 
real-time location and configuration on a PC during its operation.]

In the morning and evening, Don Pettit again sampled ppCO2 (carbon 
dioxide partial pressure) levels in the SM and Lab with the CDMK (CO2 
monitoring kit).

Budarin downlinked video recorded previously of the Russian Plasma 
Crystal-3 (PK-3) payload. He also completed his routine 24-hr. checkup 
and watering of the Russian BIO-5 Rasteniya-2/Lada-2 ("Plants-2") 
plant growth experiment.

Ken Bowersox completed the daily routine maintenance of the SOSh life
support system in the SM (incl. ASU toilet system) as well as the 
daily IMS (inventory management system) delta file preparation for 
automated export to MCC-M/TsUP. Sox also performed the daily Lab 
payload status checkup of the autonomous PCG-STES010 payload.

The planned R&R of the LAB1P5 RPCM (remote power controller module), 
which controls the VTR1 video recorder, has been deferred. [When 
trying to perform a preemptive refresh of the RPCM, the ground 
received an unexpected health flag. While it is believed that the 
event is fully understood, engineers want to assess it completely 
before proceeding.]

The crew was congratulated on their outstanding interactive TV 
downlink for PAO yesterday. Today, at 10:30am, they shared in a second 
and equally successful TV interview with a number of media clients, 
viz., CNN, ABC World News Tonight, CBS Early Show, and NBC Nightly 
News. All exchanges went exceedingly well. [As yesterday, the crew 
found confident words in responding to media questions on the topics 
of the tragic loss of Columbia, the resumption of Shuttle flights and 
the potential difficulties associated with reducing ISS crew size to 
two.]

ECLSS (environment control and life support systems) engineers on the 
ground are continuing troubleshooting (t/s) of the degraded IMV 
(intermodular ventilation). [The t/s performed by the crew last week 
was sufficient enough to focus on an examination of the flow 
straightener that is integrated on the Node IMV fan package. The 
straightener is of a honeycomb structure and approximately 
4-millimeter diameter. For the next t/s, the crew will remove the 
flexible air duct below the IMV fan and examine the flow straighteners 
with a mirror to check for blockage. Depending on the results, the 
crew may have to scavenge a corresponding Lab IMV fan for use in the 
Node location.]

Onboard water supply: There is currently a total of 986.3 L water 
onboard, which will provide approximately 164 days, assuming 
condensate recovery and processing in the Russian SRV-K system as well 
as use of potable water for technical purposes.

R&R of SM battery #7 is scheduled for tomorrow.

Today's targets for the CEO (crew earth observations program) were
Industrialized Southeastern Africa (a stagnant area of high pressure 
is located over southern Africa allowing for atmospheric aerosol 
accumulation over the region. The ISS pass was NW of most of the area 
of interest, and the crew was to try for oblique or limb views to the 
right of track to detect the depth and extent of the smog pall), Lower 
Amazon River Basin (with ISS tracking over the northwestern portion 
this target, crew was to look right of track for views of the 
southeastern estuaries of the Rios Para and Tocantins. Some sun glint 
enhancement was possible this pass), Eastern Mediterranean Dust 
(satellite imagery continues to show dust streaming northeastward off 
the Libyan-Egyptian coast. Looking obliquely right of track to 
document this phenomenon), Western Mediterranean Smog (winds and 
weather favor smog formation in central Med between the big islands of 
Sardinia and Corsica and the Italian mainland. Looking left of track 
for oblique views), and Eastern Mediterranean Smog (crew was asked to 
continue to look left of track for smog formations, especially over 
the Po River valley and the northern Adriatic Sea).

CEO images can be viewed at the website http://eol.jsc.nasa.gov

(continued)

---
* Origin: SpaceBase[tm] Vancouver Canada [3 Lines] 604-473-9357 (1:153/719)
SEEN-BY: 633/267 270
@PATH: 153/719 715 7715 140/1 106/2000 633/267

SOURCE: echomail via fidonet.ozzmosis.com

Email questions or comments to sysop@ipingthereforeiam.com
All parts of this website painstakingly hand-crafted in the U.S.A.!
IPTIA BBS/MUD/Terminal/Game Server List, © 2025 IPTIA Consulting™.